OpenStudy (anonymous):

can someone answer this quick question, i am having a brain bubble.... for 2(5x-4)^2 would you foil 5x-4 thennn multiply it by 2 or should i multiply thennn foil?

OpenStudy (anonymous):

I would personally foil first, then multiply by 2: \[2(5x-4)^{2}=2(5x-4)(5x-4)=2(25x ^{2}-20x-20x+16)=2(25x^2-40x+16)=50x^2-80x+32\]

OpenStudy (anonymous):

Whole equation did not post, but you get the point right?